OpenStudy (anonymous):

prove that sinx+cos^2x=1-2cos^2x

OpenStudy (tkhunny):

You'll need these: \(\sin^{2}(x) + \cos^{2}(x) = 1\) \(\cos(2x) = \cos^{2}(x) - \sin^{2}(x) = 1 - 2\sin^{2}(x) = 2\cos^{2}(x) - 1\)